3(3x -4) + 11 = 4x + 4
Answer by stanbon(75887)

3(3x -4) + 11 = 4x + 4
9x - 12 + 11 = 4x + 4
9x -1 = 4x + 4
5x = 5
x = 1
=============
Cheers,
Stan H.